OpenStudy (anonymous):

I really don't understand this dimension changing thing... Can someone please help me understand this? If all the dimensions of a cylinder are doubled, the volume of the cylinder will be quadrupled. True or False

OpenStudy (phi):

by dimension, they mean height, width and length. If a volume (of any shape) has each of its height, width and length doubled, the volume will go up by 2*2*2= 8 times. quadrupled means 4 times.

OpenStudy (anonymous):

oh so for this that would be true?

OpenStudy (phi):

If you multiply each dimension by 2, what do you end up multiplying the volume by ? (see above post)

OpenStudy (anonymous):

2*2=4 if it's doubled

OpenStudy (phi):

If a volume (of any shape) has each of its height, width and length doubled, the volume will go up by 2*2*2= 8 times

OpenStudy (phi):

It might be more obvious for a cube. volume of a cube is L*W*H but length, width and height are all the same number. call it L V= L*L*L if we double each dimension to 2L, the new volume is 2L*2L*2L= 8*L*L*L= 8 * old volume

OpenStudy (whpalmer4):

Another way of seeing it: imagine you have a shape that is a brick. Volume of the brick is length * width * height. \(V_{old} = l*w*h\) Now you make another brick, with each dimension being 2x that of the old brick. The formula is still the same, but here we'll have \[V_{new} = (2l)*(2w)*(2h) = 8lwh\] If we divide the new by the old, \[\frac{V_{new}}{V_{old}} = \frac{8lwh}{lwh} = 8\] So we can see that just need to plug in the scale factors of every dimension in the formula

OpenStudy (whpalmer4):

And if we only changed some of the dimensions, this works also. For example, say we made our brick twice as long, but kept the other two dimensions constant, we would have lwh = 2*1*1 = 2 so 2x the volume. I find the most useful application of this is when deciding whether or not to buy the really big pizza :-)

OpenStudy (anonymous):

Hmm @phi well for some reason my math teacher said, if it says to double a number it =the number x 2x2=4 so 4x that number... But I guess she was wrong. So it's 2*2*2=8 times. And thats all the steps? @whpalmer4 Oh I see that makes good sense... :)

OpenStudy (whpalmer4):

It's possible you misheard what your teacher said, too.

OpenStudy (anonymous):

well I typed it out in my notes what she had up on the paper... Oh well, maybe I read it wrong. :P

OpenStudy (phi):

ask your teacher again, but I think she means doubling x means 2x doubling each dimension means 2 times each: length, width and height. btw, if they ask about surface area, doubling each dimension makes the surface area of the shape by 4 times.

OpenStudy (phi):

but to finish this question, quadruple does not mean 8 times (it means 4 times)

OpenStudy (whpalmer4):

For proof that your teacher means 2x when she says doubling, give her a dollar, ask her to double it, observe whether she gives you 2 or 4 dollars in return :-)

OpenStudy (anonymous):

okay so I found the answer :) it's actually tripled

OpenStudy (anonymous):

And you both gave so great of help that I don't know who to give the medal to :P so I'll just have to fan you both! :)

OpenStudy (whpalmer4):

No, it's not. If you double all the dimensions of a cylinder, the volume goes up by 8. Volume of a cylinder with height 1 and width/length = 1 is \[V = \pi r^2h = \pi(1/2)^2(1) = \pi/4\] Volume of a cylinder with height 2 and width/length = 2 is \[V=\pi r^2h = \pi(1)^22 = 2\pi\]\[\frac{2\pi}{\pi/4} = 8\]

OpenStudy (anonymous):

oh hmm, I see... so you're saying it goes up by 8?

OpenStudy (whpalmer4):

Look, you know it has to double just from doubling the height. To only triple overall after you've doubled, means that doubling the size of the base only increases the area of the base by 1.5. That's not true — think of a square. Double the side of the square, the area goes up by 4. A circle does the same thing, in proportion.

OpenStudy (whpalmer4):

Yes, the answer is 8.

OpenStudy (anonymous):

Oh I see, I think I get it now. Thanks :)

OpenStudy (whpalmer4):

Good!